OpenStudy (anonymous):

Compute the flux of the vector field, F, through the surface, S. vector F= x i + y j + z k and S is the sphere x^2 + y^2 + z^2 = a^2 oriented outward.

OpenStudy (anonymous):

@eliassaab can you help me with this one

OpenStudy (anonymous):

@perl @dan815

OpenStudy (anonymous):

I keep getting zero

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

A parametrization of the sphere is \[(a \cos (\theta ) \sin (\phi ),a \sin (\theta ) \sin (\phi ),a \cos (\phi )),\quad 0\le \theta \le 2 \pi ; \quad 0\le \phi \le \pi\]

OpenStudy (anonymous):

The normal vector is \[\left(a^2 \cos (\theta ) \sin ^2(\phi ),a^2 \sin (\theta ) \sin ^2(\phi ),a^2 \sin (\phi ) \cos (\phi )\right)\]

OpenStudy (anonymous):

Dot multiply your field with the normal vector, you get \[a^3 \sin (\phi ) \]

OpenStudy (anonymous):

So the flux is \[\int _0^{\pi }\int _0^{2 \pi }a^3 \sin (\phi )d\theta d\phi =4 \pi a^3\]

OpenStudy (anonymous):

why is it from 0 to pi and not 0 to pi/2

OpenStudy (anonymous):

@eliassaab thank you by the way

OpenStudy (anonymous):

YW. From 0 to pi/2 you get the semi-sphere above the xy-plane